OpenStudy (anonymous):

find two numbers who sum is 64 and whose product is the greatest number possible

OpenStudy (science0229):

Hello

OpenStudy (science0229):

There is something called "Inequality of Arithmetic and Geometric Means". \[\frac{ a+b }{ 2 }\ge \sqrt{ab}\]

OpenStudy (science0229):

It states that the arithmetic mean is always greater than the geometric mean. (If a=b, then they're equal)

OpenStudy (science0229):

The information given is that a+b=64

OpenStudy (science0229):

If you solve the inequality, you get\[1024\ge ab \] So the greatest possible product is 1024.

OpenStudy (science0229):

Generally speaking, if a sum of 2 numbers are given, and you're asked to find the greatest possible product, the 2 numbers are the same. ex) The sum is 100. What is the greatest possible product of the 2 numbers? a+b=100 a=b a=50 b=50 So, the greatest possible product=ab=2500